OpenStudy (anonymous):

SAT Question: In a parallelogram. What is the value of c-b /[3a-15] --------- [b]/ | | | | /[2a-5)----------- [c]/ The answer is 30 but I don't know how to do it. I know that 2a-5 is equal to b and 3a-15 is equal to c. I know the total degree is 360. All I have is 76=a+b+c

OpenStudy (ikram002p):

graph ?

OpenStudy (anonymous):

/[3a-15] --------- [b]/ | | | | /[2a-5)----------- [c]/ Not too scale from the diagram.

OpenStudy (anonymous):

to*

OpenStudy (anonymous):

opposite angle = 180 degrees 3a-15 +c = 2a-5 +b c-b= -a +10

hartnn (hartnn):

actually opposite angles are congruent and adjacent angles sum upto 180

OpenStudy (anonymous):

oh yes xD , but same form? :o

OpenStudy (ikram002p):

first find a since 3a-15+2a-5=180

hartnn (hartnn):

3a-15= c 2a-5 =b b+c=180 3 equations 3 unknowns

OpenStudy (ikram002p):

mmm :o

OpenStudy (ikram002p):

mmm i rather that he/she find z first xD so it would be simple system

hartnn (hartnn):

add these 3a-15= c 2a-5 =b and use 3rd equation b+c =180 and you will be able to find a

OpenStudy (midhun.madhu1987):